Circular No. 6944

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cn IN NGC 3735 Michael Schwartz, Cottage Grove, OR, reports his discovery of an ap-

parent supernova (mag 15.8) on an unfiltered CCD frame taken with the 0.35-m Tenagra I automated supernova patrol telescope on June 17.48 UT. SN 1998cn is located at α = 11^{h}35^{m}54^{s}.4, δ = +70^{o}32′10′′.2 (equinox 2000.0), which is 5′′.6 west and 10′′.2 north of NGC 3735. The object is also present with unchanged brightness on a frame taken on June 18.26. The object does not appear on an unfiltered CCD image taken with the same equipment on Apr. 17.31, and the star is not present in the Vickers atlas and other CCD atlases.

SGR 1627–41 C. Kouveliotou, Universities Space Research Association (USRA); M.

Kippen, P. Woods, and G. Richardson, University of Alabama in Huntsville; V. Connaughton, National Research Council; and M. McCollough, USRA, report on behalf of the BATSE team at Marshall Space Flight Center, NASA: “BATSE has detected repeated soft γ-ray bursts consistent with the same (previously unknown) location. We recorded three bursts on June 15.109, 15.296, and 15.411 UT (BATSE triggers 6825, 6826, and 6827, respectively) with an average duration of ∼ 200 ms. Preliminary spectral analysis of the data indicates that the events are very soft, with powerlaw spectral indices varying between −3.0 and −7.0. BATSE triggered five more times, on June 17.873, 17.901, 18.004, 18.035, and 18.071 (BATSE trigger numbers 6833, 6834, 6835, 6836, and 6837, respectively). The last trigger had a peak count rate of over 300 000 counts/s above background (integrated over four detectors between 20 and 1000 keV, in 1-s intervals) and a duration of ∼ 3 s. Due to its very high peak intensity, deadtime effects prevent us from using it for localization, until we receive more data types. The weighted location of the seven remaining triggers is centered on α = 16^{h}27^{m}12^{s}, δ = −41^{o}06′.0, with an error radius of ∼ 2^{o}. This location does not correspond to any previously known soft γ-ray repeating source; we conclude that we have discovered a new source, SGR 1627–41. We have initiated an RXTE ToO and we strongly encourage wide-field observations at other wavelengths.”

1998 June 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green